OpenStudy (anonymous):

a^3+a^2b+a+b\3a^3+3a * 10a^2\2a^2-2b^2 Will someone please help me to simplify this? I evaluated it and got the following: 10a^2(a^3+a^2b+a+b) \(3a^3+3a)(2a^2-2b^2). Thanks in advanced.

OpenStudy (asnaseer):

can see how this term can be factored: 3a^3+3a

OpenStudy (asnaseer):

*can you see...

OpenStudy (anonymous):

3a(a+a^2) ??

OpenStudy (asnaseer):

not quite: 3a(a+a^2) = 3a^2 + 3a^3 which is not the same as 3a^3 + 3a

OpenStudy (asnaseer):

have another try at factoring it

OpenStudy (anonymous):

3(a^3+a)

OpenStudy (asnaseer):

good start - now notice that you can also pull out an 'a' from each term inside the brackets

OpenStudy (anonymous):

3a(a^2+a)

OpenStudy (anonymous):

i'm confusing myself...uga

OpenStudy (asnaseer):

I think you've made the same mistake as your first attempt. you correctly got to: 3a^3 + 3a = 3(a^3 + a) now, notice the terms inside the brackets both contain 'a', so we can pull out an 'a' to get: 3a^3 + 3a = 3(a^3 + a) = 3a(a^2 + 1)

OpenStudy (asnaseer):

the 'a^3' inside the bracket became 'a^2' because 'a' * 'a^2' = 'a^3' and the 'a' inside the bracket became '1' because 'a' * 1 = 'a'

OpenStudy (asnaseer):

does that make sense?

OpenStudy (anonymous):

hmm

OpenStudy (asnaseer):

if you are having trouble with this then I suggest you review this web site: http://www.coolmath.com/algebra/04-factoring/ It might help you learn more about how to factor these types of expressions

OpenStudy (anonymous):

i will take a look at it.

OpenStudy (anonymous):

when i evaluated the problem was i close at least?

OpenStudy (asnaseer):

let me know if you are still stuck after reviewing that site and I'll see if I can help further

OpenStudy (asnaseer):

you evaluated the expression correctly but did not simplify it

OpenStudy (asnaseer):

once you have learnt how to factor these expressions properly, then you will notice common terms in the numerator and denominator that can then be cancelled out to simplify the resulting expression.

OpenStudy (asnaseer):

It is worth spending some time trying to master this as it will help you a lot in future problems.

OpenStudy (asnaseer):

and, as I said, let me know afterwards if you need any further clarification - good luck! :)

OpenStudy (anonymous):

Thank you!

OpenStudy (asnaseer):

yw :)